Can someone help me solve the solution of the system using the substution method?
8x -9y = 16.5
7y - 2x = -6.5

Answer by mananth(16946)

8x-9y=16.5 .............1
7y=-6.5 +2x ---------2
/7
y=-0.93 +0.29x
Substitute y in (1)
8x-9(-0.93+0.29x )=16.5
8x+8.36 -2.57x=16.5
8x-2.57 x=16.5-8.36
5.43x=8.14
/5.43
x=1.5
Plug the value of x in (1)
8x-9y=16.5
8*1.5-9y =16.5
12-9y=16.5
-9y=4.5
/-9
y=-0.5